# Bayesian Quantile Regression with Subset Selection: A Posterior Summarization Perspective

### v2

@@ 6 Conclusion
We applied our methods to analyze the effects of environmental exposures, social stressors, and other key factors on 4th end-of-grade reading test scores for a large cohort of children in North Carolina. Our analysis revealed several important and unique insights into educational inequities. Most notably, we found that lead exposure, economic disadvantagement, racial residential isolation, and neighborhood deprivation more adversely impact reading test scores for lower-scoring students. These effects exhibited heterogeneity not only across quantiles, but also across race groups. These alarming results have important implications for childhood development and educational outcomes, and may inform policy to develop and target intervention strategies.
@@
{acks}
[Acknowledgments] The findings and conclusions in this presentation or publication are those of the authors and do not necessarily represent the views of the North Carolina Department of Health and Human Services, Division of Public Health, the National Institutes of Health, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
{funding}

### v3

@@ 6 Conclusion
We applied our methods to analyze the effects of social stressors, environmental exposures, and other key factors on 4th end-of-grade reading test scores for a large cohort of children in North Carolina. Our analysis revealed several important and unique insights into educational inequities. Most notably, we found that lead exposure, economic disadvantagement, racial residential isolation, and neighborhood deprivation more adversely impact reading test scores for lower-scoring students. These effects exhibited heterogeneity not only across quantiles, but also across race groups. Furthermore, we found that these covariates tended to be more important for prediction of the lower quantiles of educational outcomes. These alarming results have important implications for childhood development and educational attainment, and may inform policy to develop and target intervention strategies.
@@
7 Acknowledgments
The findings and conclusions in this presentation or publication are those of the authors and do not necessarily represent the views of the North Carolina Department of Health and Human Services, Division of Public Health, the National Institutes of Health, the Army Research Office, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.